A Treasury bill that settles on May 18, 2016, pays $100,000 on August 21, 2016. Assuming a discount rate of 3.23 percent, what are the price and bond equivalent yield? Use Excel to answer this question. (Round your price answer to 2 decimal places. Enter your yield answer as a percent rounded to 3 decimal places.)

Price ____

Bond equivelent yield _____
Answer #1

Bond Equivalent Yield = ((face or Par Value - Price) / Price) * (365 / D) * 100 where D is days to maturity.

Face Value or Par Value = 100,000

Purchase price or Price = 3.23% discount = Face value - (Face value x discount rate) = 100000 - (100000 x 3.23%) = 96770

Days to Maturity = difference between days (May 18,2016 and Aug 21, 2016) = 95 days

Bond Equivalent Yield = ((100000 - 96770) / 96770) *(365 / 95) * 100 = 12.824%